# FinReflectKG - EvalBench: Benchmarking Financial KG with Multi-Dimensional Evaluation

### v1

@@ Abstract
Large language models (LLMs) are increasingly being used to extract structured knowledge from unstructured financial text. Although prior studies have explored various extraction methods, there is no universal benchmark or unified evaluation framework for the construction of financial knowledge graphs (KG). We introduce FinReflectKG - EvalBench, a benchmark and evaluation framework for KG extraction from SEC 10-K filings. Building on the agentic and holistic evaluation principles of FinReflectKG - a financial KG linking audited triples to source chunks from S&P 100 filings and supporting single-pass, multi-pass, and reflection-agent-based extraction modes - EvalBench implements a deterministic commit-then-justify judging protocol with explicit bias controls, mitigating position effects, leniency, verbosity and world-knowledge reliance. Each candidate triple is evaluated with binary judgments of faithfulness, precision, and relevance, while comprehensiveness is assessed on a three-level ordinal scale (good, partial, bad) at the chunk level. Our findings suggest that, when equipped with explicit bias controls, LLM-as-Judge protocols provide a reliable and cost-efficient alternative to human annotation, while also enabling structured error analysis. Reflection-based extraction emerges as the superior approach, achieving best performance in comprehensiveness, precision, and relevance, while single-pass extraction maintains the highest faithfulness. By aggregating these complementary dimensions, FinReflectKG - EvalBench enables fine-grained benchmarking and bias-aware evaluation, advancing transparency and governance in financial AI applications.
@@
Keywords: Knowledge Graphs, LLM-as-a-Judge, Evaluation Benchmarks, SEC Filings, Large Language Models, Natural Language Processing, Financial AI
1 Background and Motivation
Large language models (LLMs) are increasingly used in finance across diverse tasks, including the extraction of information from long and unstructured documents. A common target representation for such extracted knowledge is the KnowledgeGraph (KG), which organizes information into structured triples and supports downstream applications such as compliance monitoring, risk management, and large-scale financial analytics [1, 2]. Although recent advances in LLM have improved triple extraction through multi-turn prompting and reflection [3], the evaluation of extracted triples remains a critical challenge [4]. Since financial KGs are designed to support high-stakes applications such as investment research, credit risk assessment, and portfolio decisions, the lack of reliable validation for LLM-generated triples poses significant risks.
Previous works have explored LLM-as-Judge as a scalable alternative to costly human annotation [5, 6]. However, LLM judges are susceptible to prompt sensitivity and order-based position effects [7], often favoring answers presented earlier, regardless of quality. They also tend to be overly lenient [8] and exhibit stylistic biases that reward persuasive phrasing over factual correctness [9], which undermines the robustness and interpretability of LLM-based evaluations [10, 11]. Recent studies on prompt design show that instruction-only prompting reduces reasoning drift and improves judgment consistency [12, 13, 14], whereas encouraging detailed chain-of-thought often produces the opposite effect, leading to overthinking and topic drift [15]. Moreover, evidence suggests that few-shot prompting further stabilizes LLM-as-Judge behavior and increases agreement with human annotations [16]. . Although several open-source toolkits (e.g., DeepEval, Ragas, TruLens, LangSmith) incorporate LLM-as-Judge components for tasks such as QA and summarization [17, 18, 19, 20], rigorous, bias-aware evaluation protocols tailored to triple-level KG extraction, particularly in finance, remain underspecified.
@@ 1 Background and Motivation
To address these challenges, we introduce FinReflectKG - EvalBench, a benchmark for financial KG extraction built on SEC 10-K filings. FinReflectKG - EvalBench extends the agentic principles of FinReflectKG [21] by combining schema-aware extraction with a conservative evaluation protocol. To the best of our knowledge, this is the first systematic benchmark for financial KG extraction that integrates various extraction strategies (single pass, multi pass and reflection) and a rigorous evaluation framework.
This paper makes the following key contributions:
@@
    We present FinReflectKG - EvalBench, the first benchmark for financial KG extraction from SEC filings with a reproducible and bias-aware evaluation framework.
@@
    We provide a comparative evaluation of extraction modes (single-pass, multi-pass, and reflection) across four complementary dimensions: faithfulness, precision, relevance, and comprehensiveness.
@@ 2 Methodology
We evaluate extraction outputs using an LLM-as-Judge $J$, instantiated with the Qwen3-32B model and configured for deterministic decoding (temperature =0.0=0.0). As shown in Figure 1, and following the commit-then-justify paradigm [22], the judge first produces a structured verdict and then a concise justification (up to 15 words). To further support error analysis, we introduce a warning signal that highlights extraction errors and provides actionable correction paths. These signals can also be leveraged within a feedback loop for iterative self-improvement.
@@
- Each item: {"verdict":0|1,"reasoning":"...","warning":"..."}
@@
The three extraction modes define different approaches to constructing knowledge graph triples. Single-pass uses a single LLM for both extraction and normalization. Multi-pass splits the task between two LLMs: one extracts triples, while the other normalizes them according to rubric parameters. Reflection employs an agentic, iterative workflow, where extraction and feedback loops refine triples until inconsistencies are resolved or a maximum iteration limit is reached. Across the three extraction modes, the task is to evaluate the quality of the candidate triples produced for each span $x$ with candidate set $T_{x}$. Evaluation is conducted along four complementary dimensions. First, faithfulness $F$ measures whether the content of a triple is factually grounded in the source text, without relying on world knowledge or bridging inferences. Second, precision $P$ assesses the clarity and specificity of triples, penalizing generic placeholders (e.g., “Company”) and imprecise expressions of quantities or dates. Third, relevance $R$ checks whether the triple contributes directly to the main theme of the source span rather than introducing tangential information. Finally, comprehensiveness $C$ is measured on a three-level ordinal scale (good, partial, bad) and it evaluates coverage at the chunk level, scoring how well the set of triples represents all atomic core facts.
Comprehensiveness is macro-averaged across spans:
$\bar{C}=\frac{1}{|\mathcal{X}|}\sum_{x\in\mathcal{X}}C(x).$
@@ 3 Results
Table 1: Evaluation results across extraction modes
Single Pass Multi Pass Reflection
Comprehensiveness 62.60 62.41 72.01
Faithfulness 87.25 78.73 83.40
Precision 56.06 58.01 59.49
Relevance 91.46 82.64 92.52
@@
Together, these results highlight the complementary strengths and weaknesses of different extraction strategies. Reflection provides the most balanced coverage, excelling in breadth, topical alignment, and structural accuracy, though it requires further calibration to match the strict factual reliability of single-pass generation. These findings underscore the importance of multi-dimensional evaluation, as no single metric alone captures the full spectrum of trade-offs inherent to financial KG construction.
@@ 4 Conclusions
In this work, we introduced FinReflectKG - EvalBench, the first benchmark for financial knowledge graph extraction from SEC 10-K filings with a reproducible, bias-aware evaluation framework. By integrating schema-aware extraction with a conservative LLM-as-Judge protocol, we provide a principled way to assess triple quality across multiple extraction modes. Our analysis highlights several key insights. Reflection emerges as the most balanced modes, achieving superior comprehensiveness, precision, and relevance, while single-pass remains the most faithful to the source text. This trade-off underscores the necessity of multi-dimensional evaluation, as no single metric alone can fully capture extraction quality.
Beyond reporting scores, FinReflectKG - EvalBench contributes a methodological advancement through explicit bias controls, a commit-then-justify judgment protocol, and the introduction of warning signals that enable actionable error analysis. These design elements not only improve reliability but also open the door to self-improving extraction pipelines, where diagnostic signals guide iterative refinement.
Looking forward, our benchmark provides a foundation for advancing financial KG research in several directions. Future work may extend coverage beyond the S&P 100 universe and incorporate diverse categories of financial documents. By establishing transparent, reproducible, and bias-aware evaluation standards, FinReflectKG - EvalBench aims to accelerate progress toward trustworthy financial KGs that support downstream tasks in compliance, risk management, and large-scale financial analytics.

### v2

@@ Abstract
Large language models (LLMs) are increasingly used to extract structured knowledge from unstructured financial text, yet the field lacks a standardized benchmark and unified evaluation protocol for financial knowledge graphs (KGs). We introduce FinReflectKG-EvalBench, a benchmark and bias-aware evaluation framework for KG extraction from SEC 10-K filings. Building on FinReflectKG Authors (2025), EvalBench supports single-pass, multi-pass, and reflection-based extraction, and evaluates candidate triples with a commit-then-justify LLM-as-Judge ensemble equipped with explicit bias controls (position effects, leniency, verbosity, and world-knowledge reliance). Judges assign binary labels for faithfulness, precision, and relevance, while comprehensiveness is assessed at the chunk level using a three-level scale. Inter-judge agreement is high across dimensions, supporting that the protocol yields stable and auditable judgments. Empirically, reflection achieves the strongest performance on comprehensiveness, precision, and relevance, while single-pass maintains the highest faithfulness.
@@
Keywords: Knowledge Graphs, LLM-as-Judge, Evaluation, SEC Filings, Large Language Models, Financial AI
FinReflectKG - EvalBench: Benchmarking Financial KG with Multi-Dimensional Evaluation
Fabrizio Dimino Domyn New York, US          Abhinav Arun Domyn New York, US          Bhaskarjit Sarmah Domyn Gurugram, India          Stefano Pasquali Domyn New York, US
1 Introduction
Large language models (LLMs) are increasingly used in finance to process long, complex, and unstructured documents such as regulatory filings, earnings reports, and legal disclosures. A common target representation for such extracted information is the Knowledge Graph (KG), which structures textual content into relational triples and enables applications including compliance monitoring, risk assessment, investment research, and large-scale financial analytics Li and Passino (2024); Li et al. (2024c).
Recent advances in LLM prompting strategies, such as multi-turn extraction, self-reflection, and iterative refinement, have substantially improved the quality of automatically extracted triples Zhang and Soh (2024). However, progress in extraction capability has outpaced progress in evaluation. In high-stakes financial settings, even subtle extraction errors can propagate downstream, leading to incorrect risk signals, misleading investment insights, or regulatory exposure. As a result, the lack of reliable, reproducible, and bias-aware evaluation protocols for financial KG extraction remains a critical bottleneck.
Evaluating KG extraction is inherently challenging. Financial documents are dense, context-dependent, and often require domain expertise to interpret correctly. Gold-standard annotations are expensive and difficult to scale, while exact matching metrics fail to capture semantic correctness, partial entailment, or relevance. Consequently, recent work has increasingly relied on LLM-as-Judge frameworks as a scalable alternative to human evaluation Li et al. (2024b); Evidently AI (2025). Despite their practical appeal, LLM judges introduce a new set of risks: they are sensitive to prompt phrasing and answer ordering Shi et al. (2024a), often overly lenient Ye et al. (2024), and prone to stylistic or persuasive biases that reward fluency over factual grounding Hwang et al. (2025).
These limitations are especially problematic in the context of triple-level KG extraction, where evaluation requires fine-grained judgments about faithfulness to source text, factual precision, and relevance to a predefined schema. While several open-source evaluation toolkits provide LLM-based evaluators for tasks such as question answering and summarization AI (2024); Gradients (2023); Truera (2023); LangChain (2023), they are not designed to address the unique challenges posed by structured KG extraction in regulated domains such as finance.
@@ 1 Introduction
To address this gap, we introduce FinReflectKG-EvalBench, a benchmark for evaluating financial KG extraction from SEC 10-K filings. Building on the agentic construction principles of FinReflectKG Authors (2025), our benchmark integrates schema-aware extraction with a conservative, bias-aware evaluation protocol. Unlike prior work, FinReflectKG-EvalBench systematically compares multiple extraction strategies—single-pass, multi-pass, and reflection-based extraction—under a unified evaluation framework tailored to financial risk-sensitive applications.
This paper makes the following contributions:
@@
    We introduce FinReflectKG-EvalBench, the first benchmark for financial KG extraction from SEC filings with a reproducible and bias-aware evaluation protocol.
@@
    We propose a heterogeneous ensemble of LLM judges that aggregates decisions from multiple models to mitigate prompt sensitivity, positional bias, and evaluation leniency.
  3. 3.
@@
    We present a systematic comparison of extraction strategies across four complementary dimensions: faithfulness, precision, relevance, and comprehensiveness.
  4. 4.
    We quantify evaluation stability via inter-judge agreement (Krippendorff’s $\alpha$), supporting that the protocol yields consistent judgments across dimensions.
2 Related Work
2.1 Knowledge Graph Construction from Financial Documents
Financial knowledge graphs have been constructed from a wide range of real-world artifacts, spanning regulatory filings, annual reports, brokerage research, and financial news. A recurring theme across these efforts is that financial text is both highly structured (tables, sections, standardized disclosures) and semantically ambiguous (domain jargon, entity aliases, evolving corporate structures), which makes schema design and normalization central to KG quality Li and Passino (2024); Li et al. (2024c). In practice, many pipelines therefore adopt schema, or ontology, driven representations to constrain extraction and improve interoperability, building on standardized vocabularies such as the Financial Industry Business Ontology (FIBO) and related financial ontologies EDM Council ; Bennett (2013).
Beyond filings, several works mine analyst and brokerage research reports to build domain-specific KGs that capture entities, events, and relations useful for downstream analytics. For instance, FR2KG provides a dataset and an evaluation setup for automated construction of a financial research report knowledge graph Wang et al. (2021a); while other work demonstrates scalable KG construction by mining massive brokerage research corpora, emphasizing template design, relation canonicalization, and quality control under limited supervision Cheng et al. (2022). Complementary lines of research explore knowledge discovery and graph extraction from financial texts with minimal or no labeled data, highlighting the feasibility of end-to-end pipelines but also exposing sensitivity to document heterogeneity and extraction noise Oksanen et al. (2022).
More recently, LLM-driven information extraction has expanded the design space from strictly closed schemas to more open or semi-open extraction regimes; GoLLIE shows how guideline-following instruction tuning can improve zero-shot information extraction under complex annotation schemas Sainz et al. (2024). This is particularly relevant for financial filings, where open IE over long documents can surface diverse relations that are hard to enumerate a priori; for example, recent work constructs KGs from 10-K reports using open extraction guided by relation families (e.g., business, transaction, personnel) to support downstream financial applications Guo et al. (2024). In parallel, temporal and dynamic KGs built from financial news model how entities and relations evolve over time, enabling time-aware downstream tasks such as anomaly detection and link prediction Li and Passino (2024).
As extraction pipelines become more capable, they also become more interactive and iterative. Recent prompting strategies aim to improve triple quality by allowing models to revise earlier outputs in light of schema constraints or source-text inconsistencies Papaluca et al. (2024); Mou (2024). Agentic frameworks further extend this paradigm by decomposing extraction into coordinated sub-tasks, such as entity identification, relation grounding, and consistency checking Authors (2025).
2.2 Evaluation of Knowledge Graph Extraction
Evaluating the quality of knowledge graph (KG) extraction remains an open and multifaceted challenge. Traditional metrics such as exact triple matching or surface-level overlap are brittle: they fail to account for semantic equivalence, partial correctness, or contextual relevance, and often penalize correct extractions that differ in surface form from a gold standard Chang et al. (2024). Moreover, metrics originally developed for ontology alignment or schema mapping emphasize lexical similarity and are insufficient for capturing deeper semantic relationships between extracted facts Seo et al. (2022).
Beyond matching metrics, research in KG quality assessment has proposed structural quality metrics-such as connectivity, consistency, coverage-that evaluate how well a constructed KG satisfies desirable properties such as coherence and completeness across the graph, independent of surface forms Wang et al. (2021b).
Human annotation remains the gold standard for extraction evaluation due to its ability to interpret nuanced semantics and domain-specific knowledge. However, manual annotation is costly, slow, and difficult to scale, particularly for specialized domains such as finance that require expert judgment. To address this, numerous works have employed semantic similarity and embedding-based comparisons between extracted triples and references Wei et al. (2025). These approaches use vector representations to measure semantic closeness and have been applied both for entity alignment and relation matching, but they struggle with interpretability at the schema level and may conflate distinct semantic errors.
As a result, evaluation pipelines across both information extraction and KG extraction are increasingly turning to LLM-based judges as a pragmatic compromise between scalability and semantic sensitivity. LLM-as-Judge frameworks leverage generative models to interpret and compare extracted structures in natural language, allowing for richer criteria such as faithfulness, relevance, and factuality to be incorporated into automated evaluation Li et al. (2024a). LLM judges can also produce structured critiques or rationale that assist error analysis, despite introducing their own challenges in prompt design, consistency, and calibration.
2.3 LLM-as-Judge: Reliability and Biases
A growing literature studies LLM-based evaluators not only as a convenient proxy for human assessment, but as a complex measurement instrument whose outputs depend on the evaluation protocol, prompt specification, and judge model choice. In practice, many widely-used evaluation pipelines adopt pairwise comparisons to reduce score-scale drift and improve reproducibility Zheng et al. (2023). Closely related efforts such as AlpacaEval operationalize this paradigm for instruction-following evaluation at scale, while explicitly documenting persistent confounders that can distort model rankings Li et al. (2023).
Beyond protocol choices, recent work shows that LLM-judge reliability is constrained by systematic biases that are often orthogonal to task correctness. Positional effects remain a recurring failure mode: LLM judges can favor candidates based on presentation order rather than content quality, a phenomenon documented both in benchmark-centric analyses Zheng et al. (2023) and in dedicated studies that quantify position bias across judges, tasks, and comparison settings Dimino et al. (2025); Shi et al. (2024b). Separate lines of work characterize style-related biases, where superficial attributes-such as fluency, formatting-can outweigh factual accuracy, motivating multi-dimensional rating schemes instead of collapsing evaluation into a single scalar score Wu and Aji (2023). Complementing these findings, bias taxonomies for judges also include more nuanced effects such as authority, beauty, fallacy-oversight biases under controlled perturbations, showing that even strong judges can be systematically steered Chen et al. (2024).
Importantly, judge behavior can interact with model communication strategies. For example, LLM-judges can penalize epistemic markers and expressions of uncertainty, indicating that evaluation may inadvertently disincentivize calibrated or honest uncertainty reporting Lee et al. (2025). Another recurring concern is self-preference or self-bias, where judges favor outputs that resemble their own generations or systematically assign higher scores to their own responses Chen et al. (2025). These effects suggest that judge selection and ensemble design are not merely engineering details, but central components of evaluation validity.
@@ 3.1 Corpus and span construction
We consider three extraction modes. Single-pass uses a single LLM for both extraction and normalization. Multi-pass splits the task between two LLMs: one extracts triples, while the other normalizes them according to rubric parameters. Reflection employs an agentic iterative workflow, where extraction and feedback loops refine triples until inconsistencies are resolved or a maximum iteration limit is reached.
3.2 Evaluation dimensions
Across the three extraction modes, the task is to evaluate the quality of the candidate triples produced for each span $x$ with candidate set $T_{x}$. Evaluation is conducted along four complementary dimensions. First, faithfulness $F$ measures whether the content of a triple is factually grounded in the source text, without relying on world knowledge or bridging inferences. Second, precision $P$ assesses the clarity and specificity of triples, penalizing generic placeholders (e.g., “Company”) and imprecise expressions of quantities or dates. Third, relevance $R$ checks whether the triple contributes directly to the main theme of the source span rather than introducing tangential information. Finally, comprehensiveness $C$ is measured on a three-level scale (good, partial, poor) and it evaluates coverage at the chunk level, scoring how well the set of triples represents all atomic core facts.
Comprehensiveness is macro-averaged across spans:
$\bar{C}=\frac{1}{|\mathcal{X}|}\sum_{x\in\mathcal{X}}C(x).$
3.3 LLM-as-Judge evaluation protocol
We evaluate extracted triples with a heterogeneous LLM-as-Judge ensemble $J$, instantiated with Qwen3-235B-A22B, gpt-oss-120B, and Llama-3.3-Nemotron-Super-49B-v1.5, and configured for deterministic decoding (temperature =0.0=0.0). The judges are intentionally diverse in architecture and training data, which makes agreement a practical proxy for evaluation stability: when all three judges converge on the same label, we treat the verdict as high-confidence; when they disagree, we flag the instance as ambiguous and route it to human review (human-in-the-loop).
As shown in Figure 1, and following the commit-then-justify paradigm Lopez-Lira and Tang (2023), the judge first produces a structured verdict and then a concise justification (up to 15 words). To further support error analysis, we introduce a warning signal that highlights extraction errors and provides actionable correction paths. These signals can also be leveraged within a feedback loop for iterative self-improvement.
@@ 3.3 LLM-as-Judge evaluation protocol
- Each item: {"verdict":0|1,"reasoning":"...",
"warning":"..."}
@@
3.4 Inter-judge agreement
Because our evaluation relies on a heterogeneous LLM-as-Judge panel, disagreement is not treated as noise but as a diagnostic signal: it typically concentrates on borderline cases where evidence is underspecified in the span or where multiple valid interpretations exist under strict locality constraints. To quantify stability across the three judges, we report Krippendorff’s $\alpha$ for each evaluation dimension, computed over the full set of judged instances. Unlike pairwise Cohen’s $\kappa$, $\alpha$ naturally supports multiple annotators and corrects for chance agreement while remaining interpretable under class imbalance.
Formally, for a given metric $m\in\{F,P,R,C\}$ with label space $\mathcal{Y}_{m}$, Krippendorff’s $\alpha_{m}$ is defined as
$\alpha_{m}=1-\frac{D_{o}}{D_{e}},$
where $D_{o}$ is the observed disagreement and $D_{e}$ is the disagreement expected by chance under the empirical label marginals. For comprehensiveness, we treat the tree outputs as nominal categories, since they represent discrete coverage regimes rather than a calibrated numeric scale.
@@ 4 Results
4.1 Performance across extraction modes
Table 1: Evaluation results across extraction modes.
Single Multi Refle-
Pass Pass ction
Comprehensiveness 58.75 66.32 74.27
Faithfulness 85.11 75.92 84.80
Precision 60.61 62.31 64.23
Relevance 84.35 81.39 90.44
@@ 4.1 Performance across extraction modes
4.2 Inter-judge agreement
To assess the stability of these trends under our heterogeneous LLM-as-Judge panel, we report inter-judge agreement using Krippendorff’s $\alpha$ (nominal) for each evaluation dimension, computed over the full dataset (no missing labels).
Table 2: Inter-judge agreement (Krippendorff’s $\alpha$) across evaluation dimensions.
Dimension $\boldsymbol{\alpha}$
Faithfulness 0.88
Precision 0.84
Relevance 0.77
Comprehensiveness 0.71
Agreement is highest for faithfulness ($\alpha=0.88$), consistent with the criterion largely reducing to verification of explicit textual support. Precision also exhibits strong agreement ($\alpha=0.84$), reflecting that specificity issues are comparatively easier to identify consistently. Agreement is lower for relevance ($\alpha=0.77$) and comprehensiveness ($\alpha=0.71$), which require judgments about salience and coverage and are inherently more sensitive to extraction granularity under strict locality constraints.
Overall, these agreement levels indicate that the evaluation protocol is sufficiently reliable to support comparison across extraction modes in Table 1. Taken together, the results highlight complementary strengths across strategies: reflection provides the strongest coverage and topical alignment and yields modest gains in structural accuracy, while single-pass remains the most conservative in factual grounding.
@@ 5 Conclusions
We introduced FinReflectKG-EvalBench, a benchmark for financial knowledge graph extraction from SEC 10-K filings paired with a reproducible, bias-aware evaluation framework. Our approach combines schema-aware extraction with a conservative LLM-as-Judge protocol, explicit bias controls, and a deterministic commit-then-justify procedure with actionable warning signals.
Empirically, we find a clear trade-off across extraction strategies: reflection performs best on comprehensiveness, precision, and relevance, while single-pass achieves the highest faithfulness. Inter-judge agreement is substantial across dimensions, supporting that the proposed evaluation protocol yields stable and auditable judgments for comparing extraction methods.
Overall, FinReflectKG-EvalBench enables fine-grained, multi-dimensional benchmarking of financial KG extraction and provides practical evaluation primitives that facilitate transparent error analysis and targeted human review of ambiguous cases.
Limitations
FinReflectKG-EvalBench is currently scoped to U.S. SEC Form 10-K filings from S&P 100 companies for fiscal year 2024, which limits generalization to other issuers, time periods, jurisdictions, languages, and document types (e.g., 10-Q, 8-K, earnings call transcripts, and contracts). Extending the benchmark across corpora and time is necessary to assess robustness under domain shift and temporal drift.
Our evaluation relies on an LLM-as-Judge ensemble. Although we enforce deterministic decoding, explicit bias controls, and observe substantial inter-judge agreement, LLM judges may still exhibit residual failure modes. Agreement should therefore be interpreted as protocol stability rather than ground-truth correctness; targeted human audits remain important for high-stakes deployments.
Finally, our metrics capture complementary aspects of extraction quality but do not fully characterize downstream utility. In particular, chunk-level comprehensiveness does not measure document-level coverage, cross-chunk consistency, or entity resolution over entire filings. Future work should incorporate graph consistency checks, entity linking/normalization evaluation, and task-based metrics tied to downstream financial applications.
Appendix A Appendix
Faithfulness Examples Source Text: "OpenAI signed a \$1B deal with Microsoft in 2024 in Texas" Supported (1): Triplet: ["OpenAI","Signed","\$1B deal"] Verdict: 1 (Supported), Reasoning: Triplet grounded in source text, Warning: None Not Supported (0): Triplet: ["OpenAI","Buy","Microsoft"] Verdict: 0 (Not Supported), Reasoning: Triplet not grounded in source text, Warning: Possible hallucination
Precision Examples Source Text: "OpenAI signed a \$1Billion deal with Microsoft" Precise (1): Triplet: ["OpenAI","Partners_With","Microsoft"] Verdict: 1 (Precise), Reasoning: Specific entities and relation, Warning: None Triplet: ["OpenAI","Signed","\$1Billion deal"] Verdict: 1 (Precise), Reasoning: Specific entities and amount, Warning: None Not Precise (0): Triplet: ["Company","Related_To","Something"] Verdict: 0 (Not Precise), Reasoning: Generic entity and broad relation, Warning: Generic entity Triplet: ["OpenAI","Signed","\$2Billion deal"] Verdict: 0 (Not Precise), Reasoning: Amount mismatch with text, Warning: Amount mismatch
Relevance Examples Source Text: "OpenAI signed a \$1Billion deal with Microsoft in Texas" Relevant (1): Triplet:["OpenAI","Partners_With","Microsoft"] Verdict: 1 (Relevant), Reasoning: Triplet relevant for the source text, Warning: None Not Relevant (0): Triplet: ["OpenAI","Signed","in Texas"] Verdict: 0 (Not Relevant), Reasoning: Location not relevant to the main topic, Warning: Off-topic
Comprehensiveness Examples Source Text: "In 2024, OpenAI signed a \$1Billion deal with Microsoft for AI partnership in Texas." Score 3 (Good): Set of Triplets: ["OpenAI","Partners_With","Microsoft"], ["OpenAI","Signed","\$1Billion deal"], ["OpenAI","Signed","in Texas"], ["OpenAI","Signed","for AI partnership"], ["OpenAI","Signed","in 2024"], ["Microsoft","Partners_With","OpenAI"], ["Microsoft","Signed","\$1Billion deal"], ["Microsoft","Signed","in Texas"], ["Microsoft","Signed","for AI partnership"], ["Microsoft","Signed","in 2024"] Reasoning: The set of triplets covers all core facts from the source text. Warning: None Score 2 (Partial): Set of Triplets: ["OpenAI","Partners_With","Microsoft"], ["OpenAI","Signed","\$1Billion deal"], ["OpenAI","Signed","in Texas"], ["OpenAI","Signed","for AI partnership"], ["Microsoft","Partners_With","OpenAI"], ["Microsoft","Signed","\$1Billion deal"] Reasoning: It misses the date and does not cover all for Microsoft. Warning: Possible positional bias and missing information. Score 1 (Poor): Set of Triplets: ["OpenAI","Partners_With","Microsoft"] Reasoning: It misses core facts. Warning: Incomplete set of triplets generation.